# Adaptive Cohen’s Class Time-Frequency Distribution

### v1

@@ Abstract
The fixed kernel function-based Cohen’s class time-frequency distributions (CCTFDs) allow flexibility in denoising for some specific polluted signals. Due to the limitation of fixed kernel functions, however, from the view point of filtering they fail to automatically adjust the response according to the change of signal to adapt to different signal characteristics. In this letter, we integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain. Some examples are also carried out to demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-arts in noise suppression.
Index Terms:
@@ I Introduction
From the view point of signal processing, the integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function. Namely, the CCTFD is a result of smoothing the WVD by using the kernel function. The main purpose the smoothing filter is not just only to suppress cross-terms of the WVD, but to achieve non-stationary signal denoising in the WVD domain, and accordingly extract the target signal from background noises or eliminate the impact of noises through the unique reconstruction property of WVD.
@@
There are a number of attempts to use the filtering instinct of CCTFD to remove noises, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19]. These filtering methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely. To overcome this shortcoming, it seems feasible to explore the adaptive CCTFD whose kernel function can be adjusted with the input signal automatically.
@@
Adaptive filter is a kind of filter that is able to automatically adjust its parameters or response according to the change of signal to adapt to different signal characteristics [20]. The most celebrated result in this field is Wiener’s result [21] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method. The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [22, 23, 24, 25, 26, 27, 28].
@@
The core idea of this letter is to integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to investigate the convolution type of CCTFD time-frequency analysis method. The main purpose of this letter is to design the least-squares adaptive filter method in the WVD domain, disclose the influence mechanism of the kernel function on the effect of denoising, and establish the adaptive kernel function of CCTFD with the minimum MSE in the WVD domain. The main contributions of this letter are summarized as follows:
@@
  • This letter obtains the least-squares adaptive filter in the WVD domain.
@@
  • This letter proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
@@
  • This letter demonstrates the noise suppression superiority of the proposed adaptive CCTFD over the ordinary Wiener filter and some classical fixed kernel function-based CCTFDs.
@@
The remainder of this letter is structured as follows. In Section II, we recall some necessary background and notation on the convolution type of CCTFD. In Section III, we provide the CCTFD-based adaptive filter method for additive noises jamming signals. In Section IV, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method. In Section V, we draw a conclusion. All the technical proofs of our theoretical results are relegated to the appendix parts.
@@
To facilitate the understanding and clear representation of the various symbols used throughout the paper, we present Table I, which provides a detailed explanation of each symbol and its corresponding meaning.
TABLE I: Symbol description.
@@
II Convolution type of CCTFD
@@
$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
×$\times\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
@@
Let $\ast$ and $\mathcal{F}$ be the conventional convolution operator and Fourier operator, respectively. Then, Eq. (II) can be rewritten as
@@
denotes the Fourier transform (FT) of the kernel function $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$. Eq. (2) indicates that the CCTFD is none other than the convolution of the WVD and the (FT version of) kernel function.
III Least-squares adaptive filter-based CCTFD
@@
From Eq. (II), it derives that the adaptive optimal kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain, i.e.,
@@
Substituting Eq. (13) into Eq. (II) gives the least-squares adaptive filter-based CCTFD
@@
$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}$
×$\times\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
@@
IV Numerical experiments
This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain. It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter and some classical fixed kernel function-based CCTFDs.
In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the $1$-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively. For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively.
@@ IV Numerical experiments
Figs. 1(a) and (b) plot respectively the SNR-MSE (logarithm base $10$) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using six filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, Wiener filter and the proposed adaptive CCTFD. Figs. 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these six filtering methods. Figs. 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these six filtering methods. Figs. 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these six filtering methods. It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, and Wiener filters under different SNR levels.
@@
Figure 1: Performance evaluation of six denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
@@
Figure 2: Performance evaluation of six denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
@@
Figure 3: Performance evaluation of six denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
@@
Figure 4: Performance evaluation of six denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
@@
Additionally, we investigated the denoising performance of the six methods under conditions of different colored noise. The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise. The observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4. Table II presents the MSE and PSNR values of the output signals through the six filters under the influence of three types of colored noise. It can be seen that for blue noise, our proposed adaptive CCTFD significantly outperforms the other methods. For pink and red noise, our method is superior to the others for LFM and GELFM signals (except for the PSNR value of red noise in the GELFM signal, which is slightly lower than that of the Wiener filter). For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals. However, overall, our method consistently outperforms the fixed kernel function filtering methods.
@@
TABLE II: The MSE and PSNR of six filtering methods under different colored noise conditions.
@@
Pink noise Margenau-Hill $-0.0016$ $-0.1592$ $-0.0513$ $-0.1319$
Kirkwood-Rihaczek $0.0676$ $-0.0050$ $-0.0330$ $-0.1200$
Born-Jordan $-0.1582$ $-0.3053$ $-0.1123$ $-0.2134$
Page $0.0996$ $-0.0360$ $-0.0201$ $-0.0582$
Wiener filter $-0.4985$ $-0.6911$ $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ $-0.3169$ $-0.3369$
Blue noise Margenau-Hill $0.1437$ $-0.4563$ $-0.0637$ $-0.3730$
Kirkwood-Rihaczek $0.3580$ $-0.3711$ $0.0063$ $-0.2404$
Born-Jordan $-0.1057$ $-0.5265$ $-0.2016$ $-0.9358$
Page $0.3281$ $-0.3355$ $0.0239$ $0.0682$
Wiener filter $-0.4676$ $-0.6632$ $-0.5208$ $-1.0831$
@@
Red noise Margenau-Hill $0.0655$ $-0.3004$ $-0.0624$ $-0.2233$
Kirkwood-Rihaczek $0.1243$ $-0.2567$ $-0.0475$ $-0.2145$
Born-Jordan $-0.0907$ $-0.3810$ $-0.1369$ $-0.3623$
Page $0.1720$ $-0.2790$ $-0.0082$ $-0.0851$
Wiener filter $-0.8039$ $-1.2798$ $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ $-0.4032$ $-0.3919$
@@
Pink noise Margenau-Hill $3.0323$ $5.2766$ $3.5237$ $4.3306$
Kirkwood-Rihaczek $2.3400$ $3.7410$ $3.3431$ $4.2110$
Born-Jordan $4.6190$ $7.0716$ $4.1339$ $5.1494$
Page $2.0433$ $3.7370$ $3.2109$ $3.5963$
Wiener filter $8.8314$ $10.3192$ $\mathbf{8.0556}$ $\mathbf{7.8716}$
Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ $6.1978$ $6.5799$
Blue noise Margenau-Hill $1.5865$ $7.6228$ $3.6468$ $7.2377$
Kirkwood-Rihaczek $-0.5454$ $6.8108$ $2.9469$ $5.4278$
Born-Jordan $4.1557$ $8.3757$ $5.0342$ $13.0179$
Page $-0.2493$ $6.4764$ $2.7708$ $2.3297$
Wiener filter $8.1082$ $9.8937$ $8.6899$ $14.4000$

### v2

# 2]Key Laboratory of System Control and Information Processing, Ministry of Education, Shanghai 200240, China

@@ Abstract
Inspired by the use of adaptive kernel-based Cohen’s class time-frequency distributions (CCTFDs) for cross-term suppression, this paper aims to explore novel adaptive kernel functions for denoising. We integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to design the least-squares adaptive filter method in the Wigner-Ville distribution (WVD) domain, giving birth to the least-squares adaptive filter-based CCTFD whose kernel function can be adjusted with the input signal automatically to achieve the minimum mean-square error denoising in the WVD domain. Some examples are also carried out to demonstrate that the proposed adaptive CCTFD outperforms some state-of-the-arts in noise suppression.
keywords:
@@ 1 Introduction
The integral form of CCTFD can be written as the Fourier transform (FT) of the product of the ambiguity function and the kernel function. Cross-terms may arise from the bi-linear nature of CCTFD, causing interference and reducing clarity in signal analysis. There are a number of attempts to use the kernel of CCTFD to suppress cross-terms, giving birth to many useful and effective kernel functions of which the CCTFD works well [15, 16, 17, 18, 19]. However, these methods employ fixed kernel functions in dealing with specific input signals, which limits the scope of applications severely. To overcome this shortcoming, Baraniuk et al. have proposed a series of methods using adaptive kernel functions to suppress cross-terms [20, 21, 22, 23]. The adaptive radial Gaussian kernel function-based CCTFD is one of the most famous in this field.
The integral form of CCTFD can be rewritten as the conventional convolution of the WVD and the kernel function. It is important to note that the kernel function mentioned here is a Fourier transform pair with the kernel function described above, as shown in Eq. (2). From the view point of signal processing, the CCTFD is a result of smoothing the WVD by using the kernel function. Inspired by the adaptive kernel function concept proposed by Baraniuk et al., we aim to implement adaptive denoising by designing an adaptive kernel function in the WVD domain. This approach will enable us to effectively extract the target signal from background noise and mitigate the impact of noise through the unique reconstruction property of WVD.
@@
An adaptive filter [24] is a type of filter that automatically adjusts its parameters in response to changes in the signal, making it particularly effective for noise reduction by adapting to varying characteristics of the noise and effectively suppressing it. The most celebrated result in this field is Wiener’s result [25] which applies the minimum mean-square error (MSE) criterion to design the least-squares adaptive filter method. The Wiener filter has been widely used in addressing practical issues encountered frequently in radar, communications, sonar, biomedicine, and vibration engineering [26, 27, 28, 29, 30, 31, 32].
@@
The core idea of this paper is to integrate Wiener filter principle and the time-frequency filtering mechanism of CCTFD to investigate the convolution type of CCTFD time-frequency analysis method. The main purpose of this paper is to design the least-squares adaptive filter method in the WVD domain, disclose the influence mechanism of the kernel function on the effect of denoising, and establish the adaptive kernel function of CCTFD with the minimum MSE in the WVD domain. The main contributions of this paper are summarized as follows:
@@
  • This paper obtains the least-squares adaptive filter in the WVD domain.
@@
  • This paper proposes the adaptive CCTFD whose kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain to restore signals from an extreme strong additive noise background.
@@
  • This paper demonstrates the noise suppression superiority of the proposed adaptive CCTFD over the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs, and the adaptive radial Gaussian kernel function-based CCTFD.
@@
The remainder of this paper is structured as follows. In Section 2, we recall some necessary background and notation on the convolution type of CCTFD. In Section 3, we provide the CCTFD-based adaptive filter method for additive noises jamming signals. In Section 4, we introduce some examples to validate the effectiveness, reliability and feasibility of the proposed method. In Section 5, we draw a conclusion. All the technical proofs of our theoretical results are relegated to the appendix parts.
@@
To facilitate the understanding and clear representation of the various symbols used throughout the paper, we present Table 1, which provides a detailed explanation of each symbol and its corresponding meaning.
Table 1: Symbol description.
@@
2 Convolution type of CCTFD
@@
$\mathrm{C}_{f}(\boldsymbol{x},\boldsymbol{w})=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}f\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{f\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\phi(\boldsymbol{\theta},\boldsymbol{\tau})\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta},$ (1)
@@
Let $\ast$ and $\mathcal{F}$ be the conventional convolution operator and Fourier operator, respectively. Then, Eq. (1) can be rewritten as
@@
denotes the FT of the kernel function $\phi(\boldsymbol{\theta},\boldsymbol{\tau})$. Eq. (2) indicates that the CCTFD is none other than the convolution of the WVD and the (FT version of) kernel function.
3 Least-squares adaptive filter-based CCTFD
@@
From Eq. (2), it derives that the adaptive optimal kernel function takes the reversal of the least-squares adaptive filter transfer function in the WVD domain, i.e.,
@@
Substituting Eq. (13) into Eq. (1) gives the least-squares adaptive filter-based CCTFD
@@
$=$ $\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}\int_{\mathbb{R}^{N}}g\left(\boldsymbol{y}+\frac{\boldsymbol{\tau}}{2}\right)\overline{g\left(\boldsymbol{y}-\frac{\boldsymbol{\tau}}{2}\right)}\frac{\varepsilon_{\mathrm{W}_{f},\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}{\varepsilon_{\mathrm{W}_{g}}(-\boldsymbol{\theta},-\boldsymbol{\tau})}\mathrm{e}^{-2\pi\mathrm{i}\left(\boldsymbol{\theta}\boldsymbol{x}^{\mathrm{T}}+\boldsymbol{\tau}\boldsymbol{w}^{\mathrm{T}}-\boldsymbol{y}\boldsymbol{\theta}^{\mathrm{T}}\right)}\mathrm{d}\boldsymbol{y}\mathrm{d}\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}.$ (14)
@@
4 Numerical experiments
This section performs four examples to verify the correctness and effectiveness of the least-squares adaptive filter method in the WVD domain. It also compares the denoising effect of the proposed adaptive CCTFD with that of the ordinary Wiener filter, some classical fixed kernel function-based CCTFDs and the adaptive radial Gaussian kernel function-based CCTFD.
In simulations, the fixed kernel functions are chosen as $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\phi(\boldsymbol{\theta},\boldsymbol{\tau})=\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$, (here $\left\|\cdot\right\|_{1}$ denotes the 111-norm for vectors), corresponding to the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution and the Page distribution, respectively. For simplicity, we call the filtering methods using fixed kernel functions $\cos\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)$, $\mathrm{e}^{\mathrm{i}\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$, $\frac{\sin\left(\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}\right)}{\frac{\boldsymbol{\theta}\boldsymbol{\tau}^{\mathrm{T}}}{2}}$ and $\mathrm{e}^{\mathrm{i}\boldsymbol{\theta}\left\|\boldsymbol{\tau}\right\|_{1}}$ as Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan and Page, respectively. Additionally, the filtering method utilizing the adaptive radial Gaussian kernel function-based CCTFD is referred to as the adaptive radial Gaussian kernel.
@@ 4 Numerical experiments
Figures 1(a) and (b) plot respectively the SNR-MSE (logarithm base 101010) and SNR-peak SNR (PSNR) (average of real and imaginary parts) line charts of the estimated LFM siganl using seven filtering methods including Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, Wiener filter and the proposed adaptive CCTFD. Figures 2(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate GELFM signals using these seven filtering methods. Figures 3(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate QFM signals using these seven filtering methods. Figures 4(a) and (b) plot respectively the SNR-MSE and SNR-PSNR line charts of the estimate SFM signals using these seven filtering methods. It can be seen that the proposed adaptive CCTFD achieves better noise suppression performance than Margenau-Hill, Kirkwood-Rihaczek, Born-Jordan, Page, adaptive radial Gaussian kernel, and Wiener filters under different SNR levels.
@@
Figure 1: Performance evaluation of seven denoising methods for the LFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated LFM signals; (b) PSNR performance of the estimated LFM signals.
@@
Figure 2: Performance evaluation of seven denoising methods for the GELFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated GELFM signals; (b) PSNR performance of the estimated GELFM signals.
@@
Figure 3: Performance evaluation of seven denoising methods for the QFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated QFM signals; (b) PSNR performance of the estimated QFM signals.
@@
Figure 4: Performance evaluation of seven denoising methods for the SFM signal at SNR levels from $-5\mathrm{dB}$ to $5\mathrm{dB}$ under white Gaussian noise: (a) MSE performance of the estimated SFM signals; (b) PSNR performance of the estimated SFM signals.
@@
Additionally, we investigated the denoising performance of the seven methods under conditions of different colored noise. The four signals were subjected to various types of colored noise: pink noise, blue noise, and red noise. The observation interval is set to $[-5\mathrm{s},5\mathrm{s}]$, and the sampling frequencies are set as follows: $30\mathrm{Hz}$ for example 1, $50\mathrm{Hz}$ for example 2, $150\mathrm{Hz}$ for examples 3 and 4. Table 2 presents the MSE and PSNR values of the output signals through the seven filters under the influence of three types of colored noise. It can be seen that for blue noise, our proposed adaptive CCTFD significantly outperforms the other methods. For pink and red noise, our method is superior to the others for LFM and GELFM signals (except for the PSNR value of red noise in the GELFM signal, which is slightly lower than that of the Wiener filter). For pink and red noise, the Wiener filter performs better than ours for QFM and SFM signals. However, overall, our method consistently outperforms the fixed kernel function filtering methods and the adaptive radial Gaussian kernel function filtering method.
@@
Table 2: The MSE and PSNR of seven filtering methods under different colored noise conditions.
@@
Pink noise Margenau-Hill 0.0016-0.0016- 0.0016 0.1592-0.1592- 0.1592 0.0513-0.0513- 0.0513 0.1319-0.1319- 0.1319
Kirkwood-Rihaczek 0.06760.06760.0676 0.0050-0.0050- 0.0050 0.0330-0.0330- 0.0330 0.1200-0.1200- 0.1200
Born-Jordan 0.1582-0.1582- 0.1582 0.3053-0.3053- 0.3053 0.1123-0.1123- 0.1123 0.2134-0.2134- 0.2134
Page 0.09960.09960.0996 0.0360-0.0360- 0.0360 0.0201-0.0201- 0.0201 0.0582-0.0582- 0.0582
Adaptive radial Gaussian kernel 0.0451-0.0451- 0.0451 0.4237-0.4237- 0.4237 0.0662-0.0662- 0.0662 0.0510-0.0510- 0.0510
Wiener filter 0.4985-0.4985- 0.4985 0.6911-0.6911- 0.6911 $\mathbf{-0.4699}$ $\mathbf{-0.4268}$
Adaptive CCTFD $\mathbf{-1.5318}$ $\mathbf{-0.9807}$ 0.3169-0.3169- 0.3169 0.3369-0.3369- 0.3369
Blue noise Margenau-Hill 0.14370.14370.1437 0.4563-0.4563- 0.4563 0.0637-0.0637- 0.0637 0.3730-0.3730- 0.3730
Kirkwood-Rihaczek 0.35800.35800.3580 0.3711-0.3711- 0.3711 0.00630.00630.0063 0.2404-0.2404- 0.2404
Born-Jordan 0.1057-0.1057- 0.1057 0.5265-0.5265- 0.5265 0.2016-0.2016- 0.2016 0.9358-0.9358- 0.9358
Page 0.32810.32810.3281 0.3355-0.3355- 0.3355 0.02390.02390.0239 0.06820.06820.0682
Adaptive radial Gaussian kernel 0.0481-0.0481- 0.0481 0.4892-0.4892- 0.4892 0.0671-0.0671- 0.0671 0.0191-0.0191- 0.0191
Wiener filter 0.4676-0.4676- 0.4676 0.6632-0.6632- 0.6632 0.5208-0.5208- 0.5208 1.0831-1.0831- 1.0831
@@
Red noise Margenau-Hill 0.06550.06550.0655 0.3004-0.3004- 0.3004 0.0624-0.0624- 0.0624 0.2233-0.2233- 0.2233
Kirkwood-Rihaczek 0.12430.12430.1243 0.2567-0.2567- 0.2567 0.0475-0.0475- 0.0475 0.2145-0.2145- 0.2145
Born-Jordan 0.0907-0.0907- 0.0907 0.3810-0.3810- 0.3810 0.1369-0.1369- 0.1369 0.3623-0.3623- 0.3623
Page 0.17200.17200.1720 0.2790-0.2790- 0.2790 0.0082-0.0082- 0.0082 0.0851-0.0851- 0.0851
Adaptive radial Gaussian kernel 0.01050.01050.0105 0.3455-0.3455- 0.3455 0.0743-0.0743- 0.0743 0.0395-0.0395- 0.0395
Wiener filter 0.8039-0.8039- 0.8039 1.2798-1.2798- 1.2798 $\mathbf{-0.9489}$ $\mathbf{-0.7804}$
Adaptive CCTFD $\mathbf{-1.7448}$ $\mathbf{-1.3028}$ 0.4032-0.4032- 0.4032 0.3919-0.3919- 0.3919
@@
Pink noise Margenau-Hill 3.03233.03233.0323 5.27665.27665.2766 3.52373.52373.5237 4.33064.33064.3306
Kirkwood-Rihaczek 2.34002.34002.3400 3.74103.74103.7410 3.34313.34313.3431 4.21104.21104.2110
Born-Jordan 4.61904.61904.6190 7.07167.07167.0716 4.13394.13394.1339 5.14945.14945.1494
Page 2.04332.04332.0433 3.73703.73703.7370 3.21093.21093.2109 3.59633.59633.5963
Adaptive radial Gaussian kernel 3.47583.47583.4758 7.31207.31207.3120 3.67363.67363.6736 3.52163.52163.5216
Wiener filter 8.83148.83148.8314 10.319210.319210.3192 $\mathbf{8.0556}$ $\mathbf{7.8716}$
Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ 6.19786.19786.1978 6.57996.57996.5799
Blue noise Margenau-Hill 1.58651.58651.5865 7.62287.62287.6228 3.64683.64683.6468 7.23777.23777.2377
Kirkwood-Rihaczek 0.5454-0.5454- 0.5454 6.81086.81086.8108 2.94692.94692.9469 5.42785.42785.4278
Born-Jordan 4.15574.15574.1557 8.37578.37578.3757 5.03425.03425.0342 13.017913.017913.0179
Page 0.2493-0.2493- 0.2493 6.47646.47646.4764 2.77082.77082.7708 2.32972.32972.3297
Adaptive radial Gaussian kernel 3.50463.50463.5046 7.90757.90757.9075 3.68133.68133.6813 3.20163.20163.2016
Wiener filter 8.10828.10828.1082 9.89379.89379.8937 8.68998.68998.6899 14.400014.400014.4000